graph an equation of the form Ax + By = C

x - 3y = 6
The given equation is x – 2y = 4

Answer by shantysaggu007(3)

graph an equation of the form Ax + By = C
x - 3y = 6
compare both eq.. Ax=1
By=-3
c=6
The given equation is x-3y=6
L.H.S =x-3y..........(1)
put value x=0,y=2 in eq 1
0(1)-3(2)=6
1-6=6
5=6
L.H.S ≠ R.H.S (0,2)
put value x=1,y=3 in eq 1
1(1)-3(3)=6
1-9-=6
8=6
L.H.S ≠ R.H.S (1,3)
put value x=3,y=3 in eq 1
3(1)-3(3)=6
3-9=6
6=6
L.H.S=R.H.S (3,3)
make a graph according output